OpenStudy (anonymous):

please help with these questions there are like 36 OF THEM! FREAKING OUT...

OpenStudy (anonymous):

1. Displacement is distance combined with A) Direction B) Velocity C) Speed D) Magnitude

OpenStudy (anonymous):

2. The slope of the distance-time graph is equal to the A) speed B)acceleration C) displacement D) motion

OpenStudy (anonymous):

3. Acceleration is equal to A) distance divided by time B) Change in speed divided by time C) The slope of the distance-time graph D) Change in speed multiplied by time

OpenStudy (anonymous):

Which of the following represents Newton's Third Law? A) v=d/t B)a=F/m C) F=mv D) F=0